# CausalRAG: Integrating Causal Graphs into Retrieval-Augmented Generation

### v2

@@ 5.3 Case Study
Qualitative exemplar. Figure 4 traces the end-to-end behaviour of the three systems on a single question drawn from a marketing paper: “How do different combinations of influence tactics impact the likelihood of winning a sales contract?” The semantic baseline retrieves passages that mention “influence tactics” but stop short of establishing how those tactics translate into contract wins; its answer therefore echoes the query without providing the missing mechanism. Graph-based retrieval narrows the search space by clustering entities such as buyer attention and sales contract award, yet the community summary also injects peripheral phrases, leading the model to an over-generalised explanation that only partially matches the ground truth. By contrast, the causality-driven pipeline first aligns the query with a causal pathway—Influence Tactics $\rightarrow$ Buyer Attention $\rightarrow$ Contract Award. Because that pathway is preserved through retrieval and summarisation, the generated answer pinpoints the mediating role of buyer attention and specifies which tactic pairs are most effective, fully satisfying expert judgment.
Length-controlled follow-up. To test whether these qualitative differences persist when context grows, we replicate the experiment on three versions of the same paper—abstract (250 tokens), introduction (1k), and full text (16k)—and score each system on the composite of faithfulness, precision, and recall (Figure 5). Graph-level methods benefit from additional material: their aggregate score climbs from 49.98/62.29 on the abstract to 76.37/84.27 on the full paper, confirming that community-based indexing scales gracefully with document length. Yet the causality-oriented approach remains consistently ahead, posting 72.43 on the abstract, 74.86 on the introduction, and 91.69 on the full text. This steadiness indicates that causal expansion recovers the right evidence even when local term overlap is sparse (short documents) and still filters noise when semantic matches proliferate (long documents).
Emerging pattern. Taken together, the case study and its quantitative extension suggest a general rule: graph structure raises precision by adding relational cues, but explicit causal alignment is required to preserve both precision and recall across scales. Semantic search alone misses latent mechanisms; graph aggregation alone retains residual noise. Embedding causal constraints into retrieval therefore offers a principled path to robust performance on knowledge-intensive tasks, regardless of document length or query granularity.
@@ 5.5 Conclusion and Future Work
We introduced CausalRAG, a novel framework that integrates causal reasoning into retrieval-augmented generation to address core limitations of existing RAG systems. By leveraging causality, CausalRAG retrieves context that is not only relevant but also causally grounded—enhancing generation quality, reducing hallucinations, and improving answer fidelity.
Future work may extend CausalRAG in several directions. A key area is testing scalability in long-context scenarios, which requires constructing and processing numerous high-dimensional graphs, each derived from large-scale documents spanning millions of tokens. While current practices often merge documents into a single graph or use smaller graphs due to data quality and computational constraints, this direction is critical for understanding RAG scalability. In addition to improving performance, evaluating efficiency and generation latency remains important for practical deployment.
@@
Appendix A LLM Prompts Used
The full prompts used in the CausalRAG framework for causal discovery and causal summarization (Figure 7), as illustrated in Figure 2. The prompts are adapted from Edge et al. (2024).
Appendix B OpenAlex Dataset Example
This is an exmaple (see Table 2) in the dataset Priem et al. (2022b) along with generated questions.

### v3

@@ 5.3 Case Study
Qualitative Exemplar. Figure 4 traces the end‑to‑end behaviour of the three systems on a single question drawn from a marketing paper Singh et al. (2020): “How do different combinations of influence tactics impact the likelihood of winning a sales contract?” The semantic baseline retrieves passages that mention “influence tactics” but stop short of establishing how those tactics translate into contract wins; its answer therefore echoes the query without providing the missing mechanism. Graph‑based retrieval narrows the search space by clustering entities such as buyer attention and sales contract award, yet the community summary also injects peripheral phrases, leading the model to an over‑generalised explanation that only partially matches the ground truth. By contrast, the causality‑driven pipeline first aligns the query with a causal pathway—Influence Tactics $\rightarrow$ Buyer Attention $\rightarrow$ Contract Award. Because that pathway is preserved through retrieval and summarisation, the generated answer pinpoints the mediating role of buyer attention and specifies which tactic pairs are most effective, fully satisfying expert judgment.
Length‑controlled Follow‑up. To test whether these qualitative differences persist when context grows, we replicate the experiment on three versions of the same paper—abstract (250 tokens), introduction (1k), and full text (16k)—and score each system on the composite of faithfulness, precision, and recall (Figure 5). Graph‑level methods benefit from additional material: their aggregate score climbs from 49.98/62.29 on the abstract to 76.37/84.27 on the full paper, confirming that community‑based indexing scales gracefully with document length. Yet the causality‑oriented approach remains consistently ahead, posting 72.43 on the abstract, 74.86 on the introduction, and 91.69 on the full text. This steadiness indicates that causal expansion recovers the right evidence even when local term overlap is sparse (short documents) and still filters noise when semantic matches proliferate (long documents).
Emerging Pattern. Taken together, the case study and its quantitative extension suggest a general rule: graph structure raises precision by adding relational cues, but explicit causal alignment is required to preserve both precision and recall across scales. Semantic search alone misses latent mechanisms; graph aggregation alone retains residual noise. Embedding causal constraints into retrieval therefore offers a principled path to robust performance on knowledge‑intensive tasks, regardless of document length or query granularity.
@@ 5.5 Conclusion and Future Work
We introduced CausalRAG, a novel framework that integrates causal reasoning into retrieval-augmented generation to address key limitations of existing RAG systems. Through theoretical analysis and empirical validation, we demonstrated that regular RAGs suffer from disrupted contextual integrity, semantic-over-causal retrieval, and context accuracy trade-offs. By leveraging causal graphs, CausalRAG retrieves context that is not only relevant but causally grounded—enhancing generation quality, reducing hallucinations, and improving alignment between user queries and retrieved content. Our results show consistent improvements over strong baselines across diverse domains and context lengths.
Future work can extend CausalRAG in several directions. An important future direction is to ideally evaluate scalability under long-context scenarios—requiring the construction and processing of numerous high-dimensional graphs, each derived from large-scale documents spanning millions of tokens. However, current practices often merge documents into a single graph or build smaller graphs for expert-verifiable segments, limiting real-world scalability. In addition, as LLMs evolve to support larger input sizes, developing scalable, graph-based causal retrieval methods for long-context reasoning should be an important next step.
@@
Acknowledgement
We thank the reviewers for their valuable feedback on this work. This research was supported in part by NSF Award (No. 2117439).
@@ Acknowledgement
Appendix A LLM Prompts Used
For transparency we reproduce the exact templates used in our pipeline (see Figure 7).
A.1 Causal–discovery prompt
---Role--- You are a smart assistant …
--- Goal --- Write a structured, professional causality analysis report …
---Network Data--- {graph_data}
--- Report Format --- … (see listing in the figure)
A.2 Causal–summary prompt
---Role--- You are a helpful assistant …
---Goal--- Generate a response … merge the cleaned information …
---Causal Summary--- {causal_summary}
---Target Response Length and Format--- {response_type}
---User Query--- {query}
The prompts follow three practical heuristics: (i) role conditioning to steer style; (ii) explicit section headings to simplify post-processing; (iii) a strict token budget to keep latency manageable.
@@ A.2 Causal–summary prompt
Appendix B OpenAlex Dataset Construction
We sample journal and conference papers with full text from the OpenAlex API and stratify by top-level field to avoid domain bias (see Table 1).
For each document we ask an LLM to draft five answerable questions, inspired by the procedure in Edge et al. (2024). Table 2 in this appendix shows one complete example of the resulting {document, questions} record.
@@ Appendix B OpenAlex Dataset Construction
Appendix C Implementation Details I – System Configuration
  • Base LLM. All retrieval-time and evaluation calls use gpt-4o-mini via the standard OpenAI API.
  • Embeddings. Chunks and queries are encoded with sentence-transformers all-MiniLM-L6-v2 (384 dimension, cosine similarity).
  • Vector store. Embeddings are stored in a FAISS index; retrieval depth k=3k\!=\!3 throughout.
  • Orchestration. Pipeline components are wired together with LangChain v0.2; we use its async OpenAI client to parallelise requests.
Appendix D Implementation Details II – Dataset and Evaluation
Question generation.
For every document we instruct the LLM to produce five grounded questions whose answers appear verbatim in the text. Constraining the prompts this way keeps the QA task focussed on retrieval quality rather than open-ended speculation.
Graph creation.
To assess reasoning on structured knowledge, we constructed individual graphs for each document. Although the per-document build is computationally expensive, it largely eliminates information bleed between external knowledge and gives every system the same, self-contained knowledge base.
Evaluation workload.
The experiment therefore consists of disjoint {document, question, graph} triples. For each triple we evaluate five variants running one retrieval step, one generation step, the $n$ questions, and three automatic metrics (faithfulness, context-precision, context-recall) implemented with the Ragas framework. This uniform process enables a fair, like-for-like comparison of different RAGs’ reasoning ability.